# From In Silico to In Vitro: A Comprehensive Guide to Validating Bioinformatics Findings

### v1

@@ Abstract
Index Terms:
Bioinformatics, Experimental Validation, Gene Expression, Protein-Protein Interaction, CRISPR, Next-Generation Sequencing, Artificial Intelligence, Multi-Omics, Computational Predictions
@@ I Introduction
Bioinformatics, as an interdisciplinary field, utilizes computational methods and algorithms to analyze large-scale biological data, aiming to uncover the complex patterns underlying biological processes[1, 2]. With the rapid advancement of high-throughput sequencing technologies, mass spectrometry, and big data techniques, the application of bioinformatics has expanded across various fields, from genomics to transcriptomics, proteomics to metabolomics[3, 4]. These achievements predominantly rely on computational analyses to predict gene functions, protein-protein interactions, cellular pathways, and their potential roles in diseases. However, while these computational predictions provide valuable insights, they remain theoretical hypotheses that need to be validated through experimental methods to confirm their biological relevance.
The integration of bioinformatics and wet-lab experiments plays a crucial role in biological research[5, 6]. Computational predictions are often preliminary hypotheses and require experimental validation to confirm their biological significance[7]. For instance, predictions of gene expression must be verified using quantitative PCR, RNA-Seq, or other related technologies; protein interaction networks predicted by bioinformatics need to be validated by co-immunoprecipitation, mass spectrometry, and similar techniques; and novel drug targets must be evaluated in cell-based assays and animal models. These experimental validations not only confirm the accuracy of computational predictions but also uncover new biological phenomena, advancing the respective fields of study.
Although the transition from bioinformatics predictions to experimental validation is a crucial step in biological research, this process is not straightforward. First, the selection of experimental techniques and the design of experiments must be tailored to the specific research question, as different experimental models and methods can significantly influence the results[8, 9]. For example, the choice between cell models and animal models should be based on the nature of the validation target, while different protein-protein interaction validation methods, such as Co-IP and yeast two-hybrid, come with distinct applications and technical challenges. Furthermore, experimental results may be affected by various factors such as variability in experimental conditions and the complexity of biological systems, making the correlation between computational predictions and experimental validation complex and not always directly aligned[10, 11].
@@ II Overview of Bioinformatics Methods
{forest}
[…]
@@
Bioinformatics is an interdisciplinary field that combines biology, computer science, statistics, and mathematics. It focuses on the development and application of analytical tools to handle large amounts of biological data and uncover the laws and mechanisms of life processes[43]. With the development of technologies such as genomics, transcriptomics, and proteomics, biological research has entered the ”big data era.” Bioinformatics methods not only help understand the functions of genes, proteins, and other biological molecules but also predict their interactions, providing support for disease mechanism research and drug development [44, 45]. This section will provide a detailed overview of the major types of data, methods, and tools used in bioinformatics, helping readers understand the basic framework of bioinformatics and its potential applications in experimental research [46].
@@ II-A Bioinformatics Data Types
One of the core tasks of bioinformatics is to process and analyze biological data. With the development of high-throughput sequencing technologies, the speed and scale of generating various biological data have continuously increased. Different types of data reflect different levels of biological information and provide important foundations for understanding various aspects of life activities[47]. Below are the major types of bioinformatics data.
@@ II-A1 Genomic Data
Genomic data encompasses the entire genome sequence of an organism, including coding regions, non-coding regions, regulatory regions, and other genomic elements. Through genomic technologies, such as whole-genome sequencing, researchers can obtain complete genetic information and analyze genomic features such as mutations and copy number variations (CNVs)[12]. Genomic data plays a fundamental role in biological research, from tracing evolutionary histories of species and analyzing genomic differences to discovering disease-related mutations[48]. In medical research, genomic data is widely used to identify genetic variants associated with diseases, providing data support for personalized medicine[13, 14].
@@ II-A2 Transcriptomic Data
Transcriptomic data refers to the sequence information of all RNA molecules extracted from cells. It reflects the gene expression status under specific conditions and reveals variations in transcript diversity, splicing, and transcription levels[15, 16, 17]. Transcriptomic data is commonly obtained through RNA sequencing (RNA-Seq) technologies. RNA-Seq not only allows for quantitative analysis of gene expression but also explores transcript diversity, revealing gene expression patterns under different cell types or conditions[18]. Through transcriptomic data analysis, researchers can identify differentially expressed genes (DEGs), providing new biomarkers for disease diagnosis, prognosis, and treatment strategies[49].
@@ II-A3 Proteomic Data
Proteomic data involves the types, quantities, structures, post-translational modifications (such as phosphorylation and methylation), and functions of all proteins in a cell. Proteomics research focuses not only on protein expression but also on how proteins function within the cell[20]. Through mass spectrometry (MS), researchers can perform comprehensive qualitative and quantitative analysis of cellular proteins[19]. Proteomic data analysis helps us understand cellular function networks, identify key proteins related to diseases, and even provide potential targets for drug development. For example, in cancer research, proteomics can uncover tumor-specific protein markers, advancing personalized therapy[50].
@@ II-A4 Epigenomic Data
Epigenomic data studies chemical modifications and structural changes that affect gene expression without altering the DNA sequence itself, such as DNA methylation and histone modifications[21]. Epigenomics focuses on the regulatory mechanisms of gene expression and cellular fate. Technologies such as high-throughput methylation sequencing and chromatin immunoprecipitation sequencing (ChIP-Seq) are used to obtain epigenomic data[51, 22]. Epigenomics has important applications in fields such as cancer research, aging studies, and environmental toxicology, especially since epigenomic changes are closely related to diseases like cancer and neurodegenerative diseases[52].
@@ II-B Main Methods in Bioinformatics
The core task of bioinformatics is to integrate, process, and analyze multiple types of biological data to reveal the fundamental rules and mechanisms of life. Different types of data and analytical methods provide solutions to different research questions[53]. Below are several core analytical methods in bioinformatics.
@@ II-B1 Gene Expression Analysis
Gene expression analysis aims to study how gene expression patterns change under different biological conditions [23]. Common methods for gene expression analysis include differential expression analysis, clustering analysis, and principal component analysis (PCA)[24]. Differential expression analysis is used to identify genes with significant expression changes between experimental groups (such as normal vs. disease groups); clustering analysis groups genes with similar expression patterns; PCA reduces dimensionality and identifies major expression patterns for data visualization. Gene expression analysis is crucial for understanding disease mechanisms, discovering new biomarkers, and evaluating treatment effects. For instance, differential expression analysis of transcriptomic data can identify genes related to cancer prognosis[25, 54].
@@ II-B2 Protein-Protein Interaction Prediction
Protein-protein interaction (PPI) prediction is a commonly used tool in bioinformatics to uncover the functional relationships between different proteins within the cell[26]. PPI prediction methods can be classified into several types: structure-based prediction, sequence-based similarity analysis, and network-based prediction. Structure-based prediction relies on known three-dimensional protein structures to predict potential interaction regions; sequence-based similarity analysis compares sequences in known PPI databases to predict new interactions; network-based prediction constructs PPI networks and uses graph theory algorithms to infer potential interactions[55, 27]. The construction of PPI networks provides important clues for understanding cellular functions, signaling pathways, and disease mechanisms[56].
@@ II-B3 Pathway and Network Analysis
Pathway and network analysis studies how genes, proteins, and other molecules collaborate within cells to regulate biological processes[28, 29]. By constructing gene or protein interaction networks, researchers can uncover the structure and function of complex biological pathways and networks. Common analysis methods include gene set enrichment analysis (GSEA), molecular pathway analysis, and network topology analysis. GSEA assesses the differential expression of specific gene sets under different conditions; molecular pathway analysis reveals how multiple genes or proteins interact within a particular biological process; network analysis uses graph theory algorithms to identify key nodes and network modules, further exploring the mechanisms of biological processes[30]. These analysis methods are widely used in cancer, immunology, and other fields to identify key signaling pathways associated with diseases[57].
@@ II-B4 Machine Learning and Predictive Modeling
With the complexity and large-scale growth of biological data, machine learning and artificial intelligence technologies have gradually become mainstream in bioinformatics. Machine learning can automatically learn patterns from data to build predictive models. Common machine learning methods include support vector machines (SVM)[31], random forests (RF)[58], convolutional neural networks (CNN)[32], and recurrent neural networks (RNN)[59]. Machine learning is widely applied in gene function prediction, disease diagnosis, drug screening, and other areas. For example, machine learning can predict which genes are likely to be associated with a specific disease based on gene expression data; it can also predict the potential of new drug molecules by analyzing molecular structures and activity data. The advantage of machine learning lies in its powerful data fitting and generalization ability, which allows it to handle large-scale datasets that traditional methods cannot easily address.
@@ II-C Bioinformatics Tools and Databases
To perform efficient data analysis and interpretation, bioinformatics relies on the support of various tools and databases[60]. Bioinformatics tools provide powerful data processing, analysis, and visualization capabilities, while databases offer rich resources of known biological information. Below are several common bioinformatics tools and databases.
@@ II-C1 Genome and Transcriptome Analysis Tools
Tools such as GATK (Genome Analysis Toolkit) and STAR (Spliced Transcripts Alignment to a Reference) are extensively used in bioinformatics for tasks such as quality control, alignment, variant detection, and gene expression analysis of genomic and transcriptomic data. GATK is one of the most widely adopted tools for genomic analysis, particularly in the detection of genetic variants, and it is renowned for its robustness and scalability in analyzing large datasets [33, 34]. STAR, on the other hand, is a highly efficient RNA-Seq alignment tool that is widely used for transcriptomic data analysis, offering high-speed and accurate alignment of RNA-Seq reads to a reference genome, which is essential for accurate gene expression quantification [35].
@@ II-C2 Proteomics Tools
Proteomics tools, such as MaxQuant and Proteome Discoverer, are primarily used for the analysis of mass spectrometry data, enabling both qualitative and quantitative assessment of proteins. MaxQuant is a powerful, open-source software widely used for processing and analyzing high-throughput mass spectrometry data. It supports various input data formats and provides detailed protein quantification results, making it a standard tool in the proteomics community [36, 37]. Proteome Discoverer, a commercial software, offers comprehensive mass spectrometry data analysis capabilities, integrating multiple search engines to improve the reliability and sensitivity of peptide and protein identification [61, 62].
@@ II-C3 Biological Databases
Biological databases are central to bioinformatics, enabling researchers to manage, query, and analyze vast amounts of biological data. These databases provide researchers with essential access to genomic, transcriptomic, and proteomic information. Notable biological databases include NCBI, Ensembl, Gene Ontology (GO), and Reactome. NCBI offers a wealth of genomic, transcriptomic, and proteomic data, and is one of the most comprehensive resources available for molecular biology research [38]. Ensembl provides cross-species genome annotation information, making it a valuable resource for comparative genomics [39]. Gene Ontology offers a standardized system for categorizing gene functions across different species, helping to elucidate the biological roles of genes [40]. Reactome, a curated database of biological pathways, provides detailed information on cellular signaling pathways involving genes and proteins, and is particularly useful for pathway analysis and network modeling [41].
@@ III From Computational Predictions to Wet Lab Validation
{forest}
[…]
@@
The transition from computational predictions to wet lab validation is a critical step in the process of confirming the biological relevance and experimental feasibility of bioinformatics findings. This process involves designing and performing experiments to test hypotheses or validate computational models [63]. The aim is to confirm whether the computational predictions hold true in biological systems and whether the insights derived from computational analyses can be effectively translated into experimental settings. This section will outline the general process of validation, selection of appropriate experimental models, and methods used to confirm bioinformatics predictions.

### v2

@@ I Introduction
Bioinformatics, as an interdisciplinary field, utilizes computational methods and algorithms to analyze large-scale biological data, aiming to uncover the complex patterns underlying biological processes[kasabov2005computational, benton1996bioinformatics]. With the rapid advancement of high-throughput sequencing technologies, mass spectrometry, and big data techniques, the application of bioinformatics has expanded across various fields, from genomics to transcriptomics, proteomics to metabolomics[premchand2024role, olushola2024bioinformatics]. These achievements predominantly rely on computational analyses to predict gene functions, protein-protein interactions, cellular pathways, and their potential roles in diseases. However, while these computational predictions provide valuable insights, they remain theoretical hypotheses that need to be validated through experimental methods to confirm their biological relevance.
The integration of bioinformatics and wet-lab experiments plays a crucial role in biological research[xu2022, vignani2019]. Computational predictions are often preliminary hypotheses and require experimental validation to confirm their biological significance[yehudi2022]. For instance, predictions of gene expression must be verified using quantitative PCR, RNA-Seq, or other related technologies; protein interaction networks predicted by bioinformatics need to be validated by co-immunoprecipitation, mass spectrometry, and similar techniques; and novel drug targets must be evaluated in cell-based assays and animal models. These experimental validations not only confirm the accuracy of computational predictions but also uncover new biological phenomena, advancing the respective fields of study.
Although the transition from bioinformatics predictions to experimental validation is a crucial step in biological research, this process is not straightforward. First, the selection of experimental techniques and the design of experiments must be tailored to the specific research question, as different experimental models and methods can significantly influence the results[baldi2001bioinformatics, szabo2016detecting]. For example, the choice between cell models and animal models should be based on the nature of the validation target, while different protein-protein interaction validation methods, such as Co-IP and yeast two-hybrid, come with distinct applications and technical challenges. Furthermore, experimental results may be affected by various factors such as variability in experimental conditions and the complexity of biological systems, making the correlation between computational predictions and experimental validation complex and not always directly aligned[zvelebil2007understanding, lacroix2003bioinformatics].
@@ II Overview of Bioinformatics Methods
Overview of Bioinformatics Methods Bioinformatics Data Types Genomic Data Whole-genome sequencing, Mutation analysis, Personalized medicine Transcriptomic Data RNA-Seq, Gene expression, DEGs Proteomic Data Mass spectrometry, Post-translational modifications Epigenomic Data DNA methylation, Histone modification Main Bioinformatics Methods Gene Expression Analysis Differential expression, PCA, Clustering Protein-Protein Interaction (PPI) Prediction Structure-based, Sequence-based, Network-based Pathway and Network Analysis GSEA, Network topology Machine Learning and Predictive Modeling SVM, CNN, Random Forest Bioinformatics Tools and Databases Genome and Transcriptome Tools GATK, STAR Proteomics Tools MaxQuant, Proteome Discoverer Biological Databases NCBI, Ensembl, GO, Reactome [burke2007personalized, ginsburg2009genomic, ahmed2020human][hoque2022differential, zhao2021identifying, xia2017bioinformaticsidentification, rosati2024differential][meissner2022emerging, he2003proteomics][cacabelos2015epigenetics, landgrave2015epigenetic][brazma2000gene, dervaeux2010how, vanguilder2008twenty][ghosh2024matpip, tang2023machine][wang2023co, wu2014pathway, liang2023spatial][Christmann2008, wu2017introduction][McKenna2010, VanDerAuwera2013, Dobin2013][Cox2008, huang2024seaop][wheeler2007database, Flicek2014, gene2012gene, JoshiTope2005, pasche2011whole]
@@
Bioinformatics is an interdisciplinary field that combines biology, computer science, statistics, and mathematics. It focuses on the development and application of analytical tools to handle large amounts of biological data and uncover the laws and mechanisms of life processes[mathai2020validation]. With the development of technologies such as genomics, transcriptomics, and proteomics, biological research has entered the ”big data era.” Bioinformatics methods not only help understand the functions of genes, proteins, and other biological molecules but also predict their interactions, providing support for disease mechanism research and drug development [jovic2022single, yang2017scalability]. This section will provide a detailed overview of the major types of data, methods, and tools used in bioinformatics, helping readers understand the basic framework of bioinformatics and its potential applications in experimental research [chen2017integrated].
@@ II-A Bioinformatics Data Types
One of the core tasks of bioinformatics is to process and analyze biological data. With the development of high-throughput sequencing technologies, the speed and scale of generating various biological data have continuously increased. Different types of data reflect different levels of biological information and provide important foundations for understanding various aspects of life activities[andrade_bioinformatics_1997]. Below are the major types of bioinformatics data.
@@ II-A1 Genomic Data
Genomic data encompasses the entire genome sequence of an organism, including coding regions, non-coding regions, regulatory regions, and other genomic elements. Through genomic technologies, such as whole-genome sequencing, researchers can obtain complete genetic information and analyze genomic features such as mutations and copy number variations (CNVs)[burke2007personalized]. Genomic data plays a fundamental role in biological research, from tracing evolutionary histories of species and analyzing genomic differences to discovering disease-related mutations[he2017big]. In medical research, genomic data is widely used to identify genetic variants associated with diseases, providing data support for personalized medicine[ginsburg2009genomic, ahmed2020human].
@@ II-A2 Transcriptomic Data
Transcriptomic data refers to the sequence information of all RNA molecules extracted from cells. It reflects the gene expression status under specific conditions and reveals variations in transcript diversity, splicing, and transcription levels[hoque2022differential, zhao2021identifying, xia2017bioinformaticsidentification]. Transcriptomic data is commonly obtained through RNA sequencing (RNA-Seq) technologies. RNA-Seq not only allows for quantitative analysis of gene expression but also explores transcript diversity, revealing gene expression patterns under different cell types or conditions[rosati2024differential]. Through transcriptomic data analysis, researchers can identify differentially expressed genes (DEGs), providing new biomarkers for disease diagnosis, prognosis, and treatment strategies[yang2020high].
@@ II-A3 Proteomic Data
Proteomic data involves the types, quantities, structures, post-translational modifications (such as phosphorylation and methylation), and functions of all proteins in a cell. Proteomics research focuses not only on protein expression but also on how proteins function within the cell[he2003proteomics]. Through mass spectrometry (MS), researchers can perform comprehensive qualitative and quantitative analysis of cellular proteins[meissner2022emerging]. Proteomic data analysis helps us understand cellular function networks, identify key proteins related to diseases, and even provide potential targets for drug development. For example, in cancer research, proteomics can uncover tumor-specific protein markers, advancing personalized therapy[chautard2009interaction].
@@ II-A4 Epigenomic Data
Epigenomic data studies chemical modifications and structural changes that affect gene expression without altering the DNA sequence itself, such as DNA methylation and histone modifications[cacabelos2015epigenetics]. Epigenomics focuses on the regulatory mechanisms of gene expression and cellular fate. Technologies such as high-throughput methylation sequencing and chromatin immunoprecipitation sequencing (ChIP-Seq) are used to obtain epigenomic data[cacabelos2019epigenetics, landgrave2015epigenetic]. Epigenomics has important applications in fields such as cancer research, aging studies, and environmental toxicology, especially since epigenomic changes are closely related to diseases like cancer and neurodegenerative diseases[holland2017future].
@@ II-B Main Methods in Bioinformatics
The core task of bioinformatics is to integrate, process, and analyze multiple types of biological data to reveal the fundamental rules and mechanisms of life. Different types of data and analytical methods provide solutions to different research questions[oulas2019systems]. Below are several core analytical methods in bioinformatics.
@@ II-B1 Gene Expression Analysis
Gene expression analysis aims to study how gene expression patterns change under different biological conditions [brazma2000gene]. Common methods for gene expression analysis include differential expression analysis, clustering analysis, and principal component analysis (PCA)[dervaeux2010how]. Differential expression analysis is used to identify genes with significant expression changes between experimental groups (such as normal vs. disease groups); clustering analysis groups genes with similar expression patterns; PCA reduces dimensionality and identifies major expression patterns for data visualization. Gene expression analysis is crucial for understanding disease mechanisms, discovering new biomarkers, and evaluating treatment effects. For instance, differential expression analysis of transcriptomic data can identify genes related to cancer prognosis[vanguilder2008twenty, loven2012revisiting].
@@ II-B2 Protein-Protein Interaction Prediction
Protein-protein interaction (PPI) prediction is a commonly used tool in bioinformatics to uncover the functional relationships between different proteins within the cell[ghosh2024matpip]. PPI prediction methods can be classified into several types: structure-based prediction, sequence-based similarity analysis, and network-based prediction. Structure-based prediction relies on known three-dimensional protein structures to predict potential interaction regions; sequence-based similarity analysis compares sequences in known PPI databases to predict new interactions; network-based prediction constructs PPI networks and uses graph theory algorithms to infer potential interactions[wu2024mape-ppi, tang2023machine]. The construction of PPI networks provides important clues for understanding cellular functions, signaling pathways, and disease mechanisms[bernett2024cracking].
@@ II-B3 Pathway and Network Analysis
Pathway and network analysis studies how genes, proteins, and other molecules collaborate within cells to regulate biological processes[wang2023co, wu2014pathway]. By constructing gene or protein interaction networks, researchers can uncover the structure and function of complex biological pathways and networks. Common analysis methods include gene set enrichment analysis (GSEA), molecular pathway analysis, and network topology analysis. GSEA assesses the differential expression of specific gene sets under different conditions; molecular pathway analysis reveals how multiple genes or proteins interact within a particular biological process; network analysis uses graph theory algorithms to identify key nodes and network modules, further exploring the mechanisms of biological processes[liang2023spatial]. These analysis methods are widely used in cancer, immunology, and other fields to identify key signaling pathways associated with diseases[ozpak2024neuroprotective].
@@ II-B4 Machine Learning and Predictive Modeling
With the complexity and large-scale growth of biological data, machine learning and artificial intelligence technologies have gradually become mainstream in bioinformatics. Machine learning can automatically learn patterns from data to build predictive models. Common machine learning methods include support vector machines (SVM)[Christmann2008], random forests (RF)[mila2024random], convolutional neural networks (CNN)[wu2017introduction], and recurrent neural networks (RNN)[Mienye2024]. Machine learning is widely applied in gene function prediction, disease diagnosis, drug screening, and other areas. For example, machine learning can predict which genes are likely to be associated with a specific disease based on gene expression data; it can also predict the potential of new drug molecules by analyzing molecular structures and activity data. The advantage of machine learning lies in its powerful data fitting and generalization ability, which allows it to handle large-scale datasets that traditional methods cannot easily address.
@@ II-C Bioinformatics Tools and Databases
To perform efficient data analysis and interpretation, bioinformatics relies on the support of various tools and databases[matsuoka2024bioinformatics]. Bioinformatics tools provide powerful data processing, analysis, and visualization capabilities, while databases offer rich resources of known biological information. Below are several common bioinformatics tools and databases.
@@ II-C1 Genome and Transcriptome Analysis Tools
Tools such as GATK (Genome Analysis Toolkit) and STAR (Spliced Transcripts Alignment to a Reference) are extensively used in bioinformatics for tasks such as quality control, alignment, variant detection, and gene expression analysis of genomic and transcriptomic data. GATK is one of the most widely adopted tools for genomic analysis, particularly in the detection of genetic variants, and it is renowned for its robustness and scalability in analyzing large datasets [McKenna2010, VanDerAuwera2013]. STAR, on the other hand, is a highly efficient RNA-Seq alignment tool that is widely used for transcriptomic data analysis, offering high-speed and accurate alignment of RNA-Seq reads to a reference genome, which is essential for accurate gene expression quantification [Dobin2013].
@@ II-C2 Proteomics Tools
Proteomics tools, such as MaxQuant and Proteome Discoverer, are primarily used for the analysis of mass spectrometry data, enabling both qualitative and quantitative assessment of proteins. MaxQuant is a powerful, open-source software widely used for processing and analyzing high-throughput mass spectrometry data. It supports various input data formats and provides detailed protein quantification results, making it a standard tool in the proteomics community [Cox2008, huang2024seaop]. Proteome Discoverer, a commercial software, offers comprehensive mass spectrometry data analysis capabilities, integrating multiple search engines to improve the reliability and sensitivity of peptide and protein identification [Perkins1999, orsburn2021proteome].
@@ II-C3 Biological Databases
Biological databases are central to bioinformatics, enabling researchers to manage, query, and analyze vast amounts of biological data. These databases provide researchers with essential access to genomic, transcriptomic, and proteomic information. Notable biological databases include NCBI, Ensembl, Gene Ontology (GO), and Reactome. NCBI offers a wealth of genomic, transcriptomic, and proteomic data, and is one of the most comprehensive resources available for molecular biology research [wheeler2007database]. Ensembl provides cross-species genome annotation information, making it a valuable resource for comparative genomics [Flicek2014]. Gene Ontology offers a standardized system for categorizing gene functions across different species, helping to elucidate the biological roles of genes [gene2012gene]. Reactome, a curated database of biological pathways, provides detailed information on cellular signaling pathways involving genes and proteins, and is particularly useful for pathway analysis and network modeling [JoshiTope2005].
@@ III From Computational Predictions to Wet Lab Validation
From Computational Predictions to Wet Lab Validation Overview of the Validation Process Identifying Hypotheses Experimental Model Selection Testing Predictions Refining Models Selection of Experimental Models Cell Culture Models Gene expression, Protein localization, Cellular responses Animal Models Genetic modifications, Drug testing, Disease modeling Organoid or 3D Culture Models Tissue architecture, Cellular interactions, Disease modeling [azuaje2017computational][Chen2018][baranczewski2006introduction][zhang2016review][de2020research][park2021applications][walpole2013multiscale]
@@
The transition from computational predictions to wet lab validation is a critical step in the process of confirming the biological relevance and experimental feasibility of bioinformatics findings. This process involves designing and performing experiments to test hypotheses or validate computational models [azuaje2017computational]. The aim is to confirm whether the computational predictions hold true in biological systems and whether the insights derived from computational analyses can be effectively translated into experimental settings. This section will outline the general process of validation, selection of appropriate experimental models, and methods used to confirm bioinformatics predictions.